# Tiling Spaces and the Expanding Universe: Bridging Quantum Mechanics and Cosmology

### v1

@@ 3.2 Implications of the Expanding Universe as a Growing Quasicrystal
1. Elimination of the Inflation Scenario: The necessity for an inflationary period may be obviated. The large number of tiles or points, $N$, required for uniform distribution is achieved by construction through the imposition of a quasicrystalline structure. This uniformity, intrinsic to quasicrystals, provides a natural explanation for the observed large-scale homogeneity of the universe.
2. Radiation Epoch: Following the initial QECC construction phase, the universe enters a radiation-dominated epoch characterized by the emergence of phasons and phonons. These quasiparticles are analogous to relativistic matter and radiation, contributing to the early universe’s dynamics. The presence of these quasiparticles can lead to unique dispersion relations influencing the energy density and expansion rate.
@@
3. Phason Contribution at Later Times: At later times, as the number of points $N$ increases, phasons do not significantly contribute to the energy density. This is due to the dominance of a diffusive imaginary term in their dispersion relation at large wavelengths, effectively damping their influence on the universe’s expansion.
@@
4. Matter and Long Wavelength Phonons: Matter, along with long-wavelength massive phonons, continues to play a critical role in the universe’s evolution. These components contribute to the overall energy density and can interact with ordinary matter, potentially acting as candidates for dark matter.
@@
5. Accelerated Expansion: The universe’s accelerated expansion can be explained by the internal rescaling mechanism driven by a fixed matter scale. This rescaling process results from the continuous projection and replacement of points within the quasicrystalline structure, analogous to the cosmological constant or dark energy driving the observed acceleration in the universe’s expansion.
@@
6. States and Symmetries: It is, however, questionable whether one can find exactly invariant states in a discrete space-time setting [12]. In the context of our quasicrystalline model, this suggests that the emergent large-scale symmetry is an approximation resulting from the statistical mechanics framework applied to the tiling space microstates. This leads to a partition function and energy densities that, in turn, give rise to Friedmann-like equations by imposing the Hubble law on the wavelength in the dispersion relation of propagating degrees of freedom.
@@
Overall, this model provides a novel framework for understanding the universe’s structure and dynamics, integrating concepts from condensed matter physics and cosmology to offer fresh perspectives on longstanding issues such as the Hubble tension and the nature of dark matter and dark energy.
@@
At this juncture, readers may question why projections occur, how the system determines which shells to project, and the nature of the computational framework underlying this higher dimensional lattice. They may also wonder if more lattices from higher dimensions are involved. These inquiries parallel conventional physics questions about the Big Bang, conservation laws, and the mechanisms underlying physical laws. Despite these profound questions, the hypothesis of a higher-dimensional framework introduces testable consequences for the cosmos—namely, the emergence of a quasicrystalline structure distinct from both continuous and traditional crystalline constructs.
@@
If the cosmos possesses a quasicrystalline structure, this feature could potentially be identified using techniques similar to those employed in materials science for detecting crystalline arrangements. The structural nature of a material—whether crystalline, quasicrystalline, or amorphous—is revealed through distinctive diffraction patterns. In the lab, such patterns are commonly produced through electron diffraction, where an electron beam interacts with the material. The resultant diffraction pattern, which provides detailed information on atomic spacing and structure, is then recorded. The presence of definitive Bragg peaks within these patterns indicates constructive interference from regular atomic planes, revealing the material’s symmetry and structure.
@@
Applying this concept to a cosmic scale, one might consider the cosmic microwave background (CMB) radiation as a macroscopic “diffraction pattern” reflecting the universe’s structure. The CMB, a relic radiation from the Big Bang, offers a primordial snapshot of the universe from 380,000 years after its inception, containing valuable structural and compositional data from that era. Detailed examination of temperature and polarization fluctuations within the CMB, as captured by observatories like the Planck satellite, allows researchers to piece together the universe’s large-scale structure.
@@
In a theoretical model positing quasicrystalline symmetry in the universe, this analysis would involve searching the CMB data for Bragg-like peaks—specific intensity points revealing an ordered pattern characteristic of quasicrystalline structures. Identifying such patterns would profoundly transform our understanding of the cosmic expanse and bridge celestial phenomena with quantum-scale observations in condensed matter physics.
4 Discussion
In this paper, we have expanded the interpretation of the wave function beyond its traditional role in estimating a particle’s position probability. By exploring the Schrödinger equation for a particle in a box with time-dependent boundaries and introducing a constraint that bridges microscale and macroscale quantities, we have derived an equation that bears a structural resemblance to the Friedmann equation. This development could be instrumental in addressing pivotal cosmological challenges, such as the Hubble tension. This tension arises due to observed discrepancies in the measurement of the universe’s expansion rate, as denoted by the Hubble constant $H_{0}$. Notably, $H_{0}$ shows differing values when measured in the early universe—particularly via the CMB, indicating a slower expansion rate—compared to measurements from the late universe using supernovae and Cepheid variables, which suggest a faster expansion rate. Such variations point to a potential gap in our current understanding of the universe’s expansion trajectory.
@@ 4 Discussion
Furthermore, we have hypothesized a novel heuristic model of the expanding universe conceptualized as a quasicrystalline projection from a higher-dimensional lattice. The inherent properties of quasicrystals, which resonate with cosmological principles, offer promising insights into the dynamics of the cosmic scale parameter. Our approach introduces a fresh perspective on comprehending the universe’s structure and its expansion, with a specific focus on unraveling the Hubble tension, dark energy, and dark matter. In our model, the universe’s scale parameter is envisioned as a function of both the non-linear increase in points from projections from a higher-dimensional lattice and the rescaling of spaces inside the matter content of the universe with new points coming to the projected tiling space. The model provides a phase pre-inflation with quantum error correction, bypassing the inflation phase, and transitioning to a phase dominated by propagating phasons, phonons, and radiation. Subsequently, it evolves to a phase characterized by the emergence of matter in the limit of very large number of tiles, culminating in a later phase with two kinds of expansion engines due to the projections. Long-wavelength massive phonons arise as candidates for dark matter.
@@
A potential critique of this model could be the apparent contradiction between the continuous symmetry observed in the universe and the discrete nature of a quasicrystalline structure. This critique can be addressed by noting that quasicrystals give rise to tiling spaces that exhibit continuous symmetry on very large scales. We introduced a statistical mechanics framework to show how this can occur, using a probabilistic measure for tiling space microstates, which leads to a partition function and energy densities. The Friedmann-like equations arise by imposing the Hubble law on the wavelength in the dispersion relation of propagating degrees of freedom.
To fully realize the potential of this approach, further detailed analysis is required. This will involve rigorously developing the quantitative model and enhancing its predictive power, thereby contributing significantly to our understanding of the universe’s expansion and its underlying code structure.

### v2

@@ 4.1 Implications of the Universe as a Growing Quasicrystal
    Alternative to Inflation: The need for an inflationary period may be reconsidered. In our model, the large number of tiles or points $N$ required for a uniform distribution is achieved inherently through the quasicrystalline structure. This intrinsic uniformity of quasicrystals could provide a natural explanation for the observed large-scale homogeneity and isotropy of the universe without invoking an inflationary phase.
  2. 2.
    Phasons as Scalar Fields and Reheating: Following the initial quantum growth phase, the universe transitions to an epoch characterized by the emergence of phasons and phonons—quasiparticles associated with the quasicrystalline structure. Phasons, in particular, can be modeled as scalar fields similar to the inflaton in conventional cosmology. Their dynamics can contribute to reheating processes, generating matter and radiation in the early universe. The unique dispersion relations of these quasiparticles influence the energy density and expansion rate during this period.
  3. 3.
@@
    Diminishing Phason Influence at Later Times: As the universe continues to grow and the number of points $N$ increases, the contribution of phasons to the energy density becomes negligible. This is due to the dominance of a diffusive imaginary term in their dispersion relation at large wavelengths, effectively damping their influence on the universe’s expansion in later epochs.
  4. 4.
@@
    Role of Matter and Long-Wavelength Phonons: Matter, along with long-wavelength massive phonons, remains critical in the universe’s evolution. These components contribute significantly to the overall energy density and can interact with ordinary matter. Long-wavelength massive phonons, in particular, could act as candidates for dark matter, influencing structure formation and cosmic dynamics.
  5. 5.
@@
    Accelerated Expansion and Internal Rescaling: The universe’s accelerated expansion can be explained by an internal rescaling mechanism driven by a fixed matter scale. This rescaling results from the continuous projection and replacement of points within the quasicrystalline structure. It effectively acts as a form of dark energy or cosmological constant, driving the observed acceleration in the universe’s expansion without requiring additional exotic energy components.
  6. 6.
@@
    Emergent Symmetries from Discrete Structures: While exact invariant states in a discrete spacetime setting may be challenging to identify [12], our quasicrystalline model suggests that large-scale symmetries emerge as approximations. By applying a statistical mechanics framework to the tiling space microstates, we can derive partition functions and energy densities. These lead to Friedmann-like equations when we impose the Hubble-Lemaître’s law on the wavelength in the dispersion relations of the propagating degrees of freedom. Thus, continuous symmetries can emerge from the underlying discrete quasicrystalline structure at macroscopic scales.
@@
Overall, this model provides a novel framework for understanding the universe’s structure and dynamics, integrating concepts from condensed matter physics and cosmology. It offers fresh perspectives on longstanding issues such as the Hubble tension and the nature of dark matter and dark energy.
4.2 Testing the Quasicrystalline Universe
@@
At this juncture, readers may question why projections occur, how the system determines which shells to project, and the nature of the computational framework underlying this higher-dimensional lattice. They may also wonder if higher-dimensional lattices are involved beyond the initial construction. These inquiries parallel conventional physics questions about the Big Bang, conservation laws, and the fundamental mechanisms underlying physical laws. Despite these profound questions, the hypothesis of a higher-dimensional framework introduces testable consequences for the cosmos—namely, the emergence of a quasicrystalline structure distinct from both continuous and traditional crystalline constructs.
@@
If the cosmos possesses a quasicrystalline structure, this feature could potentially be identified using techniques similar to those employed in materials science for detecting crystalline arrangements. The structural nature of a material—whether crystalline, quasicrystalline, or amorphous—is revealed through distinctive diffraction patterns. In the laboratory, such patterns are commonly produced through electron diffraction, where an electron beam interacts with the material. The resultant diffraction pattern provides detailed information on atomic spacing and structure. The presence of definitive Bragg peaks within these patterns indicates constructive interference from regular atomic planes, revealing the material’s symmetry and structure.
@@
Applying this concept on a cosmic scale, one might consider the CMB radiation as a macroscopic “diffraction pattern” reflecting the universe’s structure. The CMB, a relic radiation from the early universe, offers a primordial snapshot from approximately 380,000 years after the Big Bang, containing valuable structural and compositional data. Detailed examinations of temperature and polarization fluctuations within the CMB, as captured by observatories like the Planck satellite, allow researchers to infer the universe’s large-scale structure.
@@
In a theoretical model positing quasicrystalline symmetry in the universe, this analysis would involve searching the CMB data for Bragg-like peaks—specific intensity patterns revealing an ordered structure characteristic of quasicrystals. Identifying such patterns would have profound implications, potentially transforming our understanding of the cosmos and bridging celestial phenomena with quantum-scale observations in condensed matter physics.
4.3 Quantum Error-Correcting Codes
The growing Hilbert space structure presented earlier suggests that the early universe may be viewed as a form of a quasicrystalline topological quantum computer, with quantum error-correcting codes (QECCs) inherently embedded [34]. At each growth step, a subspace $\mathcal{C}_{n}\subset\mathcal{H}_{n}$ can be identified, spanned by a set of wavefunctions ${|\Psi{n,[T]}\rangle}$ that exhibit local indistinguishability and recoverability properties [34]:
$|\Psi_{n,[T]}\rangle=\sum_{g\in G_{n}}\beta(gT_{n})|gT_{n}\rangle,$ (42)
This growing QECC structure provides a framework for understanding how quantum information can be encoded and protected within the geometry of a pre-inflation quasicrystalline universe. Notably, the Hilbert spaces associated with certain quasicrystals are isomorphic to the fusion Hilbert spaces that appear in topological quantum computing systems [35]. This establishes a connection between QECCs and topological quantum computation within the context of our working hypothesis of the growing quasicrystalline universe.
4.4 Conclusion
In this paper, we have proposed a novel heuristic model of the expanding universe conceptualized as a quasicrystalline projection from a higher-dimensional lattice. By presenting the growing quasicrystal model first, we have established a foundational framework where the intrinsic growth dynamics of a quasicrystal can naturally give rise to the observed cosmic expansion. The inherent properties of quasicrystals, which resonate with cosmological principles such as self-similarity and aperiodic order, offer promising insights into the dynamics of the cosmic scale factor.
@@ 4.4 Conclusion
Our model envisions the universe’s scale factor as a function of both the non-linear increase in points resulting from projections from a higher-dimensional lattice and the rescaling of spaces within the matter content of the universe as new points are added to the projected tiling space. This dual growth mechanism provides an early phase characterized by quantum error correction, potentially bypassing the need for an inflationary phase. The universe then transitions to a phase dominated by propagating phasons, phonons, and radiation, eventually evolving to a matter-dominated phase in the limit of a very large number of tiles. In later epochs, the universe experiences two kinds of expansion mechanisms due to the continuous projections and internal rescaling. Long-wavelength massive phonons emerge as candidates for dark matter, offering a fresh perspective on its nature and role in cosmic evolution.
Motivated by the intrinsic quasicrystal growth, we explored the Schrödinger equation for a particle in both flat and curved spaces. By considering a time-dependent potential arising from the quasicrystal’s expansion, we introduced a constraint that bridges microscale quantum phenomena with macroscale cosmological quantities. This approach led to the derivation of an equation bearing a structural resemblance to the Friedmann equation, which governs the dynamics of the universe’s expansion in General Relativity.
This development could be instrumental in addressing pivotal cosmological challenges, such as the Hubble tension. The Hubble tension arises due to observed discrepancies in the measurement of the universe’s expansion rate, as denoted by the Hubble constant H0. Notably, H0 shows differing values when measured in the early universe—particularly via the CMB, indicating a slower expansion rate—compared to measurements from the late universe using supernovae and Cepheid variables, which suggest a faster expansion rate. Such variations point to potential gaps in our current understanding of the universe’s expansion history. Our model offers a new avenue to reconcile these differences by providing a mechanism where the expansion rate naturally varies due to the intrinsic properties of quasicrystal growth.
@@
A potential critique of this model could be the apparent contradiction between the continuous symmetries observed in the universe and the discrete nature of a quasicrystalline structure. However, this can be addressed by noting that quasicrystals give rise to tiling spaces that exhibit continuous symmetries on very large scales. We introduced a statistical mechanics framework to demonstrate how continuous behavior emerges from the underlying discrete structure. By applying a probabilistic measure to the tiling space microstates, we derived partition functions and energy densities. The Friedmann-like equations arose by imposing the Hubble-Lemaître’s law on the wavelengths in the dispersion relations of propagating degrees of freedom, showing that large-scale cosmological dynamics can be consistent with a fundamentally discrete spacetime.
Furthermore, the exploration of the Schrödinger equation within this framework underscores the interplay between quantum mechanics and cosmology. It suggests that quantum states are influenced by the expansion of spacetime inherent in the quasicrystal growth, potentially leading to observable consequences in the behavior of quantum particles. This connection between microscale and macroscale phenomena enriches our understanding of the universe and highlights the importance of considering novel models that integrate concepts from different areas of physics.
To fully realize the potential of this approach, further detailed analysis is required. Future work should focus on rigorously developing the quantitative aspects of the model, exploring its predictive power, and examining its compatibility with observational data. By advancing this research, we aim to contribute significantly to our understanding of the universe’s expansion and its underlying structure. The integration of quasicrystal mathematics with quantum mechanics and cosmology offers a rich and promising framework that could shed light on some of the most profound mysteries in physics.